What is the axis of symmetry and vertex of the graph of the function.
y=4x^2+7

Answer by nerdybill(7384)

y=4x^2+7
.
y=4x^2+0x+7
.
Axis of symmetry:
x = -b/2a = -0/(2*4) = -0/8 = 0
.
To find vertex, plug the above back into the equation and solve for y:
y=4x^2+7
y=4(0)^2+7
y=0+7
y = 7
vertex is at (0,7)
.
Answers:
axis of symmetry:
x = 0
vertex:
(0,7)